OpenStudy (anonymous):

Use rounding to the nearest whole number and clustering to give an estimate for: 15.23 + 12.8 + 4. 93 + 7.21 40 41 42

OpenStudy (anonymous):

what is clustering?

OpenStudy (anonymous):

15.23 + 12.8 + 4. 93 + 7.21 15 + 12 + 4 + 7 = 27 + 11 = 38 .23 + .8 + .93 + .21 = 2.17 38 + 2.17 = ?